OpenStudy (anonymous):

integral of (dx)/(4-9x) I got -(1/9)*log(4-9x)+C But, my homework program keeps saying it's wrong... someone help?

OpenStudy (anonymous):

your homework program is mistaken stick to your answer

OpenStudy (anonymous):

maybe it is a syntax problem maybe i wants \(\ln(4-9x)\) rather than \(\log(4-9x)\)

OpenStudy (anonymous):

I tried that.. It says evaluate the indefinite integral integral (dx)/(4-9x)

zepdrix (zepdrix):

Maybe it wants the Natural bars | | for the contents of the log.

OpenStudy (anonymous):

the what bars?..

OpenStudy (anonymous):

@zepdrix good suggestion

OpenStudy (anonymous):

\[-\frac{1}{9}\log(|4-9x|)+C\] try

OpenStudy (anonymous):

RIGHT! :) THANK YOU

zepdrix (zepdrix):

Oh cool :) yay team!